A body is dropped from certain height and another body is dropped from same height after 2 secs. What will be the separation between two bodies 4 second after the dropping of second body.

Grade:11

2 Answers

Arun
5 years ago
You can use Newton's second equation of motion here.
The first ball has traveled for 6 seconds under gravity with acceleration g and initial velocity=0.
The second ball has traveled for 4 seconds under gravity with acceleration g and initial velocity=0.
Distance between the 2 balls at the end of 6 seconds after the first ball is dropped = 1/2*9.8(36-16)
= 98m
Stuti
3 years ago
98 meters is right ans
H1=1/2×9.8×6×6
H1=176.4
Now,H2=1/2×9.8×4×4
         H2=78.4
Thus,the seperation between two bodies are 
H=H1-H2
H=178.4-78.4
H=98 meters